OpenStudy (rayford):

Solve. x2 – 10x = 5? i know it's easy but i forgot -_-

OpenStudy (wach):

You want to combine these terms into a quadratic equation, factor out, and then find x :) Solve. x2 – 10x = 5 subtract 5 from both sides: x2 - 10x - 5 = 0 Now, find two numbers that when added equal -10 and when multiplied equal -5. Can you think of any?

OpenStudy (asnaseer):

again, you can use the same methods that were explained to you by @helder_edwin in this question: http://openstudy.com/updates/501452c0e4b0fa246730cbe4

OpenStudy (wach):

In that case, what asnaseer linked applies. Good luck ~

OpenStudy (rayford):

dang it asnaseer! ur messing up my learning :{

OpenStudy (asnaseer):

how? which parts do you still not understand?

OpenStudy (rayford):

where wach left off..its hard for me to review stuff...

OpenStudy (asnaseer):

look at what helder_edwin showed you in the previous question and at least make an attempt at answering this question using those methods. start one step at a time and I'll guide you if you are going wrong.

OpenStudy (wach):

There are actually multiple ways to solve that equation. You can plug it into the formula given to you earlier or factor/complete the square. The other method is better for equations wherein the answer isn't 'whole'.

OpenStudy (rayford):

i tried solving it and i got..5+or-sqrt15. am i correct?

OpenStudy (asnaseer):

almost

OpenStudy (asnaseer):

if you show your individual steps then I can show you where you may have made a mistake

OpenStudy (asnaseer):

@wach yes there are multiple ways of solving these types of equations, but I /think/ rayford has been taught to use the quadratic equation - correct me if I'm wrong rayford.

OpenStudy (rayford):

yes u are right..

OpenStudy (asnaseer):

remember the standard form of a quadratic equation is:\[ax^2+bx+c=0\]so your first step should be to identify the values for a, b and c. what do you think they are for this question?

OpenStudy (rayford):

x2 +10x+5?

OpenStudy (asnaseer):

you don't have the signs correct - look at what wach wrote above.

OpenStudy (rayford):

x2-10x+5

OpenStudy (asnaseer):

are you looking at what wach wrote above? you still don't have all the signs correct.

hero (hero):

Hint: Add 25 to both sides

hero (hero):

x^2 - 10x + 25 = 5 + 25

OpenStudy (rayford):

thanks hero!